OpenStudy (anonymous):

How many possible positive, negative and complex solutions are there in this equation? x^3-10x^2+21x -108 = 0

OpenStudy (anonymous):

you can have; x^3-10x^2= 0, x^3+21x=0, x^3-108=0, -10x^2+21x=0, -10x^2-108=0, 21x-108=0, and then you could have the whole equation x^3-10x^2+21x-108=0

OpenStudy (anonymous):

Count the sign changes between terms.

OpenStudy (anonymous):

(Descartes' Rule of Signs)

OpenStudy (shubhamsrg):

you have x^3 - 7x^2 - 3x^2 + 21x = 108 (x-7)(x^2 - 3x) = 108 x(x-7)(x-3) = 108 you may compare x(x-7)(x-3) = 9 * 2 *6 so roots will be 9 , 2, 6..as can be predicted from descartes easily..

zepdrix (zepdrix):

 Yes, rule of signs :) Do you remember that one andrea? :D